### v3

# On Continuous Full-Order Integral-Terminal Sliding Mode Control with Unknown A Priori Bound on Uncertainty

Abstract
This study aims to solve the problem of designing a continuous and finite-time control for a class of nonlinear systems in the presence of matched uncertainty with an unknown apriori bound. First, we propose a Full-Order Integral-Terminal Sliding Manifold (FOITSM) with a conventional (discontinuous) sliding mode to show that it provides the combined attributes of both nonsingular terminal and integral sliding mode algorithms. Secondly, an Adaptive Disturbance Observer (ADO) has been designed to mitigate the impact of the uncertainty. The application of ADO-based Full-Order Integral-Terminal Sliding Mode Control (FOITSMC) substantially reduces chattering in control input under conditionally known matched disturbances. Additionally, the adaptive gains of ADO are updated non-monotonically, avoiding over-bounding of acting disturbances while ensuring global boundedness of state trajectories. Lastly, the proposed algorithm has been effectively applied to the attitude stabilization of a rigid spacecraft.
keywords:
Integral sliding mode control, Terminal sliding mode control, Adaptive disturbance observer, Chattering-free control, Adaptive sliding mode control
journal: ISA Transactions
[…]
[label1]organization=Department of Electrical Engineering, IIT Bombay ,addressline=Powai, city=Mumbai, postcode=400076, state=Maharashtra, country=India
[…]
[label2]organization=Department of Electrical Engineering, IIT Kharagpur ,addressline=Kharagpur, city=Paschim Medinipur, postcode=721302., state=West Bengal, country=India
[…]
[label3]organization=Department of Electrical Engineering, NIT Silchar ,addressline=Silchar, city=Cachar, postcode=788010, state=Assam, country=India
1 Introduction
1.1 Literature survey
Initiation of control amid uncertainties has been one of the main topics of concern for the past few decades. Sliding Mode Control (SMC) [1, 2, 3, 4, 5, 6] is one of the control strategies that converges the state trajectories to a predetermined manifold in finite time. The state trajectories are constrained to evolve in that manifold and thereby, the time evolution of state trajectories becomes insensitive to external disturbances and converges to the origin asymptotically (Conventional sliding mode [7]) or in a finite time (Terminal sliding mode [8, 9, 10, 11, 12, 13, 14]). Manifestation of such control algorithms requires a high frequency switching leading to an undesired phenomenon called ”chattering”. An SMC is generally composed of two phases; the reaching phase, the state trajectories are forced to reach the manifold from its initial condition, and the sliding phase, the state trajectories are compelled to remain invariant to that manifold. In contrast to a conventional linear sliding mode (LSM), the terminal sliding mode control (TSMC) algorithm converges states on finite time and the integral sliding mode (ISM) [15, 16, 17, 18, 19, 20, 21] is devoid of any reaching phase because it ensures the invariance of state trajectories from the initial instant, and hence, guarantees robustness from its inception. Some partial knowledge about the acting external disturbances, preferably the upper bound of the disturbances, must be known in all these algorithms. The apriori knowledge of the matched disturbances is anonymous in several practical applications or bound to differ with time. Our proposed algorithm incorporates this anonymity in the design of the control law, eliminating the reaching phase and significantly reducing chattering in the control input.
Several adaptive sliding mode control (ASMC) algorithms have been proposed to alleviate this barrier. Among many possibilities of adaptation, monotonically increasing the gain [22, 23, 24, 25] associated with the signum structure of the control algorithm may result in high gain [26]. In a recent development [27], a novel ASMC technique is proposed which doesn’t require a known apriori bound on the disturbance and ensures uniformly ultimate boundedness of state trajectories. In [28], an adaptive second-order sliding mode controller has been introduced using the Lyapunov approach, based on the assumption of boundedness of uncertainties rather than the boundedness of their time-derivative. Our proposed methodology further relaxes these assumptions by only requiring the time derivative of acting disturbances to be bounded by an unknown constant value, offering a more flexible and robust solution.
Several approaches are available in the literature regarding continuous adaptive SMC algorithms, preferably adaptive super-twisting control (STC) algorithms [29, 30]. All these approaches can broadly be classified into two main categories. The first category [26, 31, 32, 33] involves increasing the gain of the STC until the second-order sliding (2SM) is obtained; once 2SM has been obtained, adaptive gains stop evolving and thereby, remain constant. The second category [34, 35, 36] comprises the development of adaptive algorithms with minimum gain required to achieve 2SM and, thereby, ensuring that there is no unnecessary overestimation of adaptive gains while ensuring sliding. This category has some advantages over the previous one- specifically from a chattering mitigation viewpoint. To the best of the authors’ knowledge, the number of research articles from a gain minimization perspective is still considerably less and our study aims in this direction using a Lyapunov-based analysis of closed-loop stability to global uniformly ultimately boundedness (GUUB) of state trajectories and estimation errors. We propose a provably robust methodology to manage acting disturbances even when the exact bound of their time-derivative is a constant but unknown. Moreover, this proposed adaptive framework avoids over-estimating adaptive gains. We demonstrate the effectiveness of our proposed control law by comparing our results with an existing methodology and applying it to the attitude stabilization of rigid spacecraft.
1.2 Contributions
In this manuscript, our main contributions are as follows.
  • 1.
    Introducing a Full-order Integral-Terminal Sliding Manifold (FOITSM) deprived of any such reaching phase, ensuring robustness from the initial instant. Besides, in sliding phase, the state trajectories would converge in finite time.
  • 2.
    In the presence of conditionally known disturbances, the control algorithm ensures the globally uniformly boundedness of state trajectories. This conditioning relies on knowledge about the disturbance, specifically that the absolute value of its time-derivative remains bounded. However, the specific bound is unknown to us as an apriori.
  • 3.
    Implementing a disturbance observation technique to mitigate the effect of acting matched disturbances and alleviate the undesired ”chattering” phenomenon.
  • 4.
    Designing adaptive control laws without overestimation of adaptive gains.
1.3 Organisation of the paper
The manuscript unfolds as follows: In Section 2 we delve into the formulation of a layout of a novel FOITSM for a general $n^{\text{th}}$ order chain of integrators afflicted by bounded matched uncertainties with known apriori bounds. An illustrative example is presented in Section 3 for validation of the proposed manifold along with STC. To counterfeit the hurdle of unknown apriori bound on uncertainty, an ADO-based controller has been designed in Section 4. Section 5 shows an illustrative example of the ADO-based FOITSMC control of a 3rd-order chain of integrators. The performance comparison with an existing algorithm is presented in Section 6. In Section 7, an application to attitude Control of a rigid spacecraft has been manifested to show the effectiveness of the proposed disturbance-observer-based controller using the novel manifold in Section 2. Ultimately, the paper draws to a close with a concise summary of conclusions in Section 8.
1.4 Notations
Throughout this paper, $\mathbb{R}$, $\mathbb{R}^{\text{n}}$, $\mathbb{R}^{+}$ and $\mathbb{R}^{\geqslant 0}$ denote, respectively, the set of all real numbers, $n$-dimensional vectors with real components, set of all positive real numbers and nonnegative real numbers. $\mathcal{B}_{r}(x)$ denotes a ball of radius $r$ centered around $x\in\mathbb{R}^{\mathrm{n}}$. The notation $[x]^{\alpha}\text{sgn}(x)$ means $[|x_{1}|^{\alpha}\text{sign}(x_{1}),\cdots,|x_{n}|^{\alpha}\text{sign}(x_{n})]^{T}$, $\forall\alpha\in\mathbb{R},x\in\mathbb{R}^{n}$ and $\text{sgn}(y)$ implies the standard signum function with its argument as $y\in\mathbb{R}$. Further, $|\textbf{x}|=[|x_{1}|,\cdots,|x_{n}|]^{T}$ for $\mathbf{x}\in\mathbb{R}^{n}$ and a summation convention as $a_{1}^{i}\cdots a_{n}^{i}b_{i}^{1}\cdots b_{i}^{m}=\sum_{i}a_{1}^{i}\cdots a_{n}^{i}b_{1}^{i}\cdots b_{m}^{i}$ The function $||\;||:\mathcal{M}\to\mathbb{R}^{+}$ denotes the Euclidean norm on a manifold $\mathcal{M}\subset\mathbb{R}^{\mathrm{n}}$, defined as $||\mathbf{x}||=\sqrt{x^{2}_{1}+x^{2}_{2}+\cdots+x^{2}_{n}}$ for any $\mathbf{x}\in\mathcal{M}$. An open ball of radius $r$ at time $t$ is denoted as $\mathfrak{B}_{r}(\mathbf{x},t)\coloneqq\left\{\mathbf{x}\in\mathcal{M}\;\text{s.t}\;||\mathbf{x}||<r\;\;\text{at time}\;t\right\}$. For the sake of brevity, arguments of a function may be omitted where the context is quite clear.
2 Sliding mode control of nonlinear systems
Consider an $\mathrm{n}^{\mathrm{th}}$ order perturbed chain of integrators which is affine in control
$\left\{\begin{array}[]{l}\dot{x}_{1}=x_{2}\\ \dot{x}_{2}=x_{3}\\ \vdots\\ \dot{x}_{n}=f(\textbf{x},t)+b(\textbf{x},t)u(t)+d_{0}\left(\textbf{x},t\right)\end{array}\right.$ (5)
where $\textbf{x}(t)=\left[x_{1}(t),x_{2}(t),\cdots,x_{n}(t)\right]^{T}\in\mathbb{R}^{\mathrm{n}}$ comprises all the states in a vector representation, $f:\mathbb{R}^{\mathrm{n}}\times\mathbb{R}^{\geqslant 0}\to\mathbb{R}$ is the nonlinear drift which is considered to be locally Lipschitz, $b:\mathbb{R}^{\mathrm{n}}\times\mathbb{R}^{\geqslant 0}\to\mathbb{R}$ is the gain for control input, $u(t)\in\mathbb{R}$ is the control input and $d_{0}:\mathbb{R}^{\mathrm{n}}\times\mathbb{R}^{\geqslant 0}\to\mathbb{R}$ is the unknown perturbation representing nonparametric uncertainties and external disturbances. Further
$f(\textbf{x},t)$ $=f_{n}\left(\textbf{x},t\right)+f_{\Delta}\left(\textbf{x},t\right),$
where $f_{n}$ is the nominal component of $f$ and $f_{\Delta}$ is the perturbation from the known nominal value.
Assumption 1.
The parametric perturbations $f_{\Delta}\left(\textbf{x},t\right)$ is globally bounded, i.e. there exists some real positive constants $f_{\mathrm{max}}$ such that $|f_{\Delta}\left(\textbf{x},t\right)|\leqslant f_{\mathrm{max}}$ where $f_{\mathrm{max}}\in\mathbb{R}^{\geqslant 0}$.
Assumption 2.
The matched uncertainty $d_{0}\left(\textbf{x},t\right)$ is globally bounded, i.e. $|d_{0}\left(\textbf{x},t\right)|\leqslant d_{\mathrm{max}}$ where $d_{\mathrm{max}}\in\mathbb{R}^{\geqslant 0}$
The primary intent of the sliding mode control approach is to robustly hold on to an effectively chosen constraint, say $s$, through a high-frequency switching control. The purpose is to guide the state trajectories to precisely follow or lie in the vicinity of the chosen constraint.
Definition 2.1 (Ideal sliding [30]).
For $t_{r}\geqslant t_{0}$, if $s(t)\equiv 0,\quad\forall\;t\geqslant t_{r}$, then an ideal sliding has been obtained for system (5) on the sliding manifold $s(t)=0$.
Definition 2.2 (Real sliding [30]).
For $t_{r}\geqslant t_{0}$, $\varepsilon>0$, if $s(t)\in\mathfrak{B}_{\varepsilon}(s,t),\quad\forall\;t\geqslant t_{r}$, then an real sliding has been obtained for system (5) on the sliding manifold $s(t)=0$.
In real-time applications, control algorithms operate with discrete measurements and thereby, introduce imperfections in switching. Thus, it’s obdurate to obtain ideal sliding in such cases. However, real sliding can be obtained with a notable relaxation in sliding accuracy and effectively reducing chattering.
Consider a full-order integral terminal sliding manifold (FOITSM) of the following form
$\left\{\begin{array}[]{l}s=s_{0}(\textbf{x})-z,\\ \dot{z}=-C_{n}|x_{n}|^{\alpha_{n}}\text{sgn}(x_{n})-C_{n-1}|x_{n-1}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})\\ -...-C_{1}|x_{1}|^{\alpha_{1}}\text{sgn}(x_{1}),\end{array}\right.$ (6)
where $s_{0}(\textbf{x})=x_{n}(t)$, $z(0)=s_{0}(0)$, and $C_{i}$’s and $\alpha_{i}$’s are constants which are chosen such that the polynomial $p^{n}+C_{n}p^{n-1}+C_{n-1}p^{n-2}+\cdots+C_{2}p+C_{1}$ is Hurwitz stable and $\alpha_{i}$’s are determined based on the following conditions
$\left\{\begin{array}[]{l}\alpha_{1}=\alpha,\;\;n=1,\\ \alpha_{i-1}=\frac{\alpha_{i}\alpha_{i+1}}{2\alpha_{i+1}-\alpha_{i}},\;\;i=1,\cdots,n,\;\;\forall n\geqslant 2,\end{array}\right.$
where $\alpha_{n+1}=1$, $\alpha_{n}=\alpha$, $\alpha\in(1-\varepsilon,1)$, $\varepsilon\in(0,1)$.
Remark 1.
The condition $z(0)$ in (6) ensures that $s=0$ at initial time $t=0$. Hence, the sliding motion occurs from time $t=0$ and thus, there is no reaching phase, unlike conventional TSMC.
Remark 2.
The term “full order” refers to the sliding manifold having a dimension n for an $n^{\mathrm{th}}$ order plant, in contrast to the conventional sliding mode algorithms where the dimension of the sliding manifold is $n-1$ for an $n^{\mathrm{th}}$ order plant.
The dynamical equations (5) and (6) can be combined and represented as
$\left\{\begin{array}[]{l}\dot{x}_{1}=x_{2}\\ \dot{x}_{2}=x_{3}\\ \vdots\\ \dot{x}_{n-1}=s+z\\ \dot{z}=-C_{n}|{x_{n}}|^{\alpha_{n}}\text{sgn}(x_{n})-C_{n-1}|{x_{n-1}}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})\\ -...-C_{1}|{x_{1}}|^{\alpha_{1}}\text{sgn}(x_{1})\\ \dot{s}=f_{n}(\textbf{x},t)+b(\textbf{x},t)u(\textbf{x})+d(\textbf{x},t)+C_{n}|{x_{n}}|^{\alpha_{n}}\text{sgn}(x_{n})\\ +C_{n-1}|{x_{n-1}}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})+...+C_{1}|{x_{1}}|^{\alpha_{1}}\text{sgn}(x_{1})\end{array}\right.$ (15)
where $d\left(\textbf{x},t\right)=d_{0}\left(\textbf{x},t\right)+f_{\Delta}\left(\textbf{x},t\right)$.
Remark 3.
For a specific control $u$, if $s$ is constrained to persistently remain zero for all time $t\geqslant 0$, the dynamical variables in (15) and consequently in (5) follow a trajectory in phase space resembling the solution set of the dynamical equation
$\dot{x}_{n}=$ $-C_{n}|{x_{n}}|^{\alpha_{n}}\text{sgn}(x_{n})-C_{n-1}|{x_{n-1}}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})$
$-...-C_{1}|{x_{1}}|^{\alpha_{1}}\text{sgn}(x_{1}),$ (16)
which converges to the origin in finite time.
Theorem 2.3.
The state trajectories of the nonlinear system (15) and correspondingly trajectories of (5) under the assumptions 1 and 2, converge to the origin within finite time if the sliding manifold is chosen as in (6). The control input is determined as follows,
$u=$ $b^{-1}(\textbf{x},t)(u_{\text{eqv}}+u_{\text{dis}}),$ (17)
$u_{eqv}=$ $-f_{n}(\textbf{x},t)-C_{n}|{x_{n}}|^{\alpha_{n}}\text{sgn}(x_{n})$
$-C_{n-1}|{x_{n-1}}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})-...-C_{1}|{x_{1}}|^{\alpha_{1}}\text{sgn}(x_{1}),$ (18)
$u_{dis}=$ $-(\eta+d_{\text{max}}+f_{\text{max}})\text{sgn}(s),$ (19)
where $\eta>0$.
Remark 4.
The discontinuous control $u_{dis}$ in (19) can be replaced by a continuous STC as follows,
$\left\{\begin{array}[]{l}u_{con}=-k_{1}|s|^{\frac{1}{2}}\text{sgn}(s)+v\\ \dot{v}=-k_{2}\text{sgn}(s)\end{array}\right.$ (20)
where $k_{1}$ and $k_{2}$ are parameters defined as $k_{1}=1.5\sqrt{\rho}$ and $k_{2}=1.1\rho$ for $|\dot{d}|\leqslant\rho$.
3 Illustrative example I
To illustrate the fact stated above, let us consider a system of the form,
$\left\{\begin{array}[]{l}\dot{x}_{1}=x_{2}\\ \dot{x}_{2}=x_{3}\\ \dot{x}_{3}=u+d\end{array}\right.$ (21)
where the disturbance $d=\text{sin}(2\pi t)$ and $u$ as a control input same as (17), (18) and (20). For simulation, a sliding surface has been considered of the following form.
$s$ $=x_{3}-z,$
$\dot{z}$ $=15\;\text{sgn}(x_{3})|x_{3}|^{\frac{7}{10}}+66\;\text{sgn}(x_{2})|x_{2}|^{\frac{7}{13}}+80\;\text{sgn}(x_{1})|x_{1}|^{\frac{7}{16}}$ (22)
Following eqs (17)-(19), the corresponding input $u$ is chosen as
$u=$ $-15\;\text{sgn}(x_{3})|x_{3}|^{\frac{7}{10}}-66\;\text{sgn}(x_{2})|x_{2}|^{\frac{7}{13}}$
$-80\;\text{sgn}(x_{1})|x_{1}|^{\frac{7}{16}}+u_{con},$
where $u_{con}$ is same as eq (20) with parameters $k_{1}=3.96$ and $k_{2}=7.7$.
(a) Evolution of states on application of STC
(b) Sliding variable of FOITSMC with STC
(c) Input for FOITSMC with STC
Figure 1: 1a and 1b show the convergence of states in finite time and sliding variable remaining $0,\;\forall t\geqslant 0$, hence, demonstrating the effectiveness of FOITSM. On application of STC, a chattering-free control is shown in 1c.
All simulations are done utilising MATLAB R2021a with an ODE 4 (Runge-Kutta) solver assigned a fixed step size of 0.001. Results are illustrated in fig. (1a), (1b) and (1c). In figure (1a), all state trajectories distinctly converge to the origin within finite time. Fig (22) shows the sliding variable starting and persisting at zero, indicating the absence of a reaching phase.. A continuous control $u$, devoid of any singularity as well as chattering (shown in Fig. (1c)), has been applied.
4 Chattering free adaptive SMC with unknown apriori bound on uncertainty
For the plant (21), consider a sliding dynamics $s$ as in eq. (6) with the nature of the disturbance $d\in C^{1}$ as
$|\dot{d}(t)|\leqslant k,$ (23)
where $k$ is a constant but unknown. The choice of $u$ should be such that $s$ becomes zero in a finite time.
Theorem 4.1.
The system (5), characterized by disturbance $d$ as defined in (23), is globally uniformly stable (GUB) and globally uniformly ultimately bounded (GUUB) ([37]) in finite time $t_{r}$
$t_{r}$ $\leqslant\frac{1}{\gamma}\text{ln}\left[\frac{V(0)-\frac{\bar{\delta}}{\gamma}}{\bar{\delta}\left(\frac{1}{\gamma-\theta}-\frac{1}{\gamma}\right)}\right],$ (24)
if the control $u$ is chosen as
$u=$ $b^{-1}(\textbf{x},t)(u_{\text{eqv}}+u_{\text{ado}}),$ (25)
$u_{\text{eqv}}=$ $-f_{n}(\textbf{x},t)-C_{n}|{x_{n}}|^{\alpha_{n}}\text{sgn}(x_{n})$
$-C_{n-1}|{x_{n-1}}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})-...-C_{1}|{x_{1}}|^{\alpha_{1}}\text{sgn}(x_{1}),$
$u_{\text{ado}}=$ $-\kappa s-\hat{d},$ (26)
and the disturbance observer dynamics
$\hat{d}=$ $\lambda\left(x_{n}-\zeta\right),$ (27)
$\dot{\zeta}=$ $f_{n}(\textbf{x},t)+b(\textbf{x},t)u+\hat{d}-\frac{\hat{k}}{\lambda}\text{sgn}(\tilde{d})-\frac{s}{\lambda},$ (28)
with an ultimate bound on $||\mathcal{S}||^{2}$ as
$\mathcal{B}=\sqrt{\frac{2\bar{\delta}}{\gamma-\theta}},$
and the adaptation law
$\dot{\hat{k}}$ $=-\tau\hat{k}+\mu|s|,\;\;\hat{k}(0)>0,$ (29)
where $\mathcal{S}=\left[s,\;\tilde{d},\;\tilde{k}\right]^{T}$, $z(0)=x_{n}(0)$, $s(0)=s_{0}$, $\hat{k}(0)\in\mathbb{R}^{>0}$, $\tau=\left(\mu+1+\tau_{0}\right)$, $\lambda,\mu,\tau_{0},\kappa\in\mathbb{R}^{+}$, $\bar{\delta}=\frac{1}{2}\tau k^{2}$. The error between actual ($d$) and estimated disturbance ($\hat{d}$) is given by
$\tilde{d}=d-\hat{d},$
and, the error etween the actual $k$ and an estimation of $k$ as $\hat{k}$ is given by
$\tilde{k}$ $=k-\hat{k}.$
Remark 5.
The governing disturbance observer dynamics in (27) and (28), a dynamical equation in $\zeta$, is continuous since the discontinuity lies in $\dot{\zeta}$.
Figure 2: Schematic Diagram for Stabilisation of (5)
Fig. 2 presents a schematic diagram illustrating the proposed control algorithm. The disturbance observer in (28) reconstructs the disturbance considering the plant states, the control signal, the adaptive gains and the sliding variable and feeding them to the controller. The controller in (26) uses a state feedback mechanism also making use of the sliding variable, estimated disturbance and the adaptive gain data manufactured using adaptive laws. This architecture ensures good performance in the presence of matched disturbance of bounded nature but the bounds being unknown.
Remark 6.
A slight modification to the control law (26) as
$u_{\text{f}_{\text{ado}}}=-\kappa_{1}s-\kappa_{2}|s|^{\frac{1}{2}}\text{sgn}(s),$
for $\kappa_{1}>\frac{1}{2}\mu$, $\kappa_{2}>0$ would result in a Fast FOITSMC with a faster convergence rate.
Remark 7.
The Lyapunov stability analysis of the closed-loop system uniquely incorporates the dynamics of disturbance error, which sets it apart from other designs of disturbance observers found in existing literature.
Remark 8.
In eq (28), a term containing $\text{sgn}(\tilde{d})$ has been used. Since $d$ is not available for measurement, we define a function,
$\omega(t)=$ $\int_{t_{0}}^{t}\tilde{d}(\sigma)\mathrm{d}\sigma,$
$=$ $x_{n}-\int_{t_{0}}^{t}\left(f(\mathbb{x},\sigma)+b(\mathbb{x},\sigma)u(\sigma)+\hat{d}(\mathbb{x},\sigma)\right)\mathrm{d}\sigma,$
and define $\text{sgn}(\tilde{d})$ as $\text{sgn}(\tilde{d}(t))=\text{sgn}\left(w(t)-w(t-\tau_{d})\right)$ where $\tau_{d}$ is the time delay.
Since $\tilde{d}=\underset{t\to 0}{\text{lim}}\frac{\left(w(t)-w(t-\tau_{d})\right)}{\tau_{d}}$, one can choose $\tau_{d}$ as fundamental sampling time of the system. It is to be noted we need to know the sign of $\tilde{d}$ instead of $\tilde{d}$ whose calculation is comparatively easy.
Remark 9.
The dynamical equation (28) constitutes the disturbance observer ([38, 39]) for the system (5) with the characterisation of disturbance $d$ as given in (23).
5 Illustrative Example II
Consider a $3^{\text{rd}}$ order chain of integrator as
$\left\{\begin{array}[]{l}\dot{x}_{1}=x_{2}\\ \dot{x}_{2}=x_{3}\\ \dot{x}_{3}=u+d\end{array}\right.$ (30)
where the disturbance $d=\text{sin}(2\pi t)$. The sliding surface is chosen the same as (22). The control input $u$ is designed as,
$u=$ $-15\text{sgn}(x_{3})|x_{3}|^{\frac{7}{10}}-66\text{sgn}(x_{2})|x_{2}|^{\frac{7}{13}}$
$-80\text{sgn}(x_{1})|x_{1}|^{\frac{7}{16}}+u_{ado}$ (31)
and $u_{ado}$ is designed as in (26), (28) and (29) with parameters $\kappa=5$, $\lambda=5$, $\tau=5$ and $\mu=2$.
(a) States with adaptive ADO based FOITSMC
(b) Sliding variable of adaptive ADO based FOITSMC
(c) Input for FOITSMC with adaptive ADO
(d) Disturbance estimation error for adaptive ADO based FOITSMC
(e) Estimation of $k$ for FOITSMC with adaptive ADO
Figure 3: 3a, 3b and 3d show the ultimate boundedness of state trajectories, sliding variable and disturbance estimation error, respectively, on the application of adaptive ADO-based FOITSMC. Moreover, the minimum estimated adaptive gain required for GUUB is shown in 3e.
Remark 10.
The strategies employed in adaptive STC [26, 31, 34, 35] involve signum functions in adaptation laws, which may potentially lead to unnecessary overestimated (underestimated) gains. This can occur when the gain value increases even as the magnitude of the sliding variable decreases (or increases). In contrast, the adaptive law in (29) is devoid of any such discontinuous functions and strives to attain the least possible value of $\hat{k}$ to obtain the desired result.
These simulation results are obtained using the ODE 4 (Runge-Kutta) solver of MATLAB R2021a with a fixed step size of 0.001. In fig. (3a) and fig. (3b), the states of (30) and sliding variable, respectively, are circumscribed within a specified bound in finite time on the implementation of input in eq. (31). Fig. (3c) displays a control input free from any chattering phenomenon. The estimated $k$ is non-monotonic (as shown in fig. 3e) and it decreases (or increases) depending on whether the magnitude of the sliding variable decreases (or increases).
6 Performance Comparison
To demonstrate the effectiveness of the proposed algorithm, we present a performance comparison with existing methods. For this purpose, we consider the Adaptive Super-Twisting control law (ASTW) outlined by ([40])for reference.
Theorem 6.1 ([40]).
Consider a system of the following form
$\dot{\sigma}(t)$ $=w+a_{1}(x,t)$ (32)
such that $|\dot{a}_{1}(x,t)|<\delta<\infty$, $\delta$ unknown. For any given arbitrary initial conditions $x(0)$, $\sigma(0)$, there exists a finite $t_{f}>0$ such that a real $2-$sliding mode has been established $\forall\;t>t_{f}$ via control law
$\left\{\begin{array}[]{l}\omega=-\alpha|\sigma|^{\frac{1}{2}}\text{sgn}(\sigma)+v\\ \dot{v}=-\frac{\beta}{2}\text{sgn}(\sigma)\end{array}\right.$ (33)
with adaptive gains $\alpha$, $\beta$ defined as follows
$\dot{\alpha}$ =$=\begin{cases}\omega_{1}\sqrt{\frac{\gamma_{1}}{2}}\text{sgn}\left(|\sigma|-\mu\right),\;\;&\alpha>\alpha_{m}\\ \eta,\;&\alpha\leqslant\alpha_{m}\end{cases}$ (34)
$\beta$ $=2\varepsilon\alpha$ (35)
where $\varepsilon$, $\lambda$, $\gamma_{1}$, $\omega_{1}$, $\eta$ are positive constants. The parameter $\alpha_{m}$ is an arbitrary-small constant.
(a) The adaptive gain $(\hat{k})$ for disturbance
$d=\sin(2\pi t)$
(b) The adaptive gain $(\hat{k})$ for disturbance
$d=\sin(2\pi t)+\texttt{ramp}(t)$
(c) The sliding variable $(s)$ in presence of
disturbance $d=\sin(2\pi t)$
(d) The sliding variable $(s)$ in presence of
disturbance $d=\sin(2\pi t)+\texttt{ramp}(t)$
[…]
7 Case study
The attitude control and stabilization of a rigid spacecraft in the presence of external disturbances have been a topic of interest for control enthusiasts for quite a long time ([41, 42, 43, 44, 45, 46]) because of its highly nonlinear dynamical nature which makes the control algorithm design much more challenging and persevering. This section portrays an application of the aforementioned control algorithm for the attitude stabilization of a rigid spacecraft. The unit-quaternion-based attitude dynamics of a rigid spacecraft are given by
$\begin{cases}\dot{q}_{0}&=-\frac{1}{2}q_{v}^{T}\Omega\\ \dot{{q}_{v}}&=0.5\left(q_{0}I_{3}+q_{v}^{\times}\right)\Omega\\ J\dot{\Omega}&=-\Omega^{\times}J\Omega+u+d.\end{cases}$
The unit-quaternion is defined by $q=[q_{0}\;q_{1}\;q_{2}\;q_{3}]^{T}=[q_{0}\;q_{v}]^{T}$ satisfying a constraint $q_{0}^{2}+q_{v}^{T}q_{v}=1$, where $q_{v}\in\mathbb{R}^{3}$ is the vector part and $q_{0}\in\mathbb{R}$ is the scalar component, $\omega\in\mathbb{R}^{3}$ is the angular velocity of the spacecraft, $u\in\mathbb{R}^{3}$ is the control torque, $J\in\mathbb{R}^{3\times 3}$ is the symmetric positive definite inertia matrix, $I_{3}$ is the $3\times 3$ identity matrix, and $\times$ is an operator on any vector $a=[a_{1}\;a_{2}\;a_{3}]^{T}$ such that,
$a^{\times}$ $=\begin{pmatrix}0&-a_{3}&a_{2}\\ a_{3}&0&-a_{1}\\ -a_{2}&a_{1}&0\end{pmatrix},$
Assumption 3.
The inertia matrix $J$ is assumed to be known.
Assumption 4.
The exogenous matched disturbance $d$ is continuously differentiable such that each components $d_{i}$ of $d$ satisfy,
$\dot{d}_{i}\leqslant k_{i},\;\;\text{for}\;i=1,2,3,$
where $k_{i}$s are finite.
Consider a sliding surface of the following form,
$s=e-e_{0}+\int_{t_{0}}^{t}[e]^{\frac{1}{2}}\text{sgn}(e)\mathrm{d}\tau,$
where $e=\Omega+k_{v}q_{v}$, $k_{v}>0$. The control input is chosen as follows
$u=$ $\Omega^{\times}J\Omega-\frac{1}{2}k_{v}J\left(q_{0}I_{3}-q_{v}^{\times}\right)\Omega-J[e]^{\frac{1}{2}}\text{sgn}(e)-J\Theta s$
$\hat{d}=$ $\Lambda\left(\Omega-z\right)$
$\dot{z}=$ $-J^{-1}\Omega^{\times}J\Omega+J^{-1}u+J^{-1}\hat{d}$
$-\Lambda^{-1}[\hat{k}]\text{sgn}(\tilde{d})-\Lambda^{-1}J^{-1}s$
and adaptation law as
$\dot{\hat{k}}=$ $-\tau\hat{k}+\mu|s|,\;\;\hat{k}(0)>0,$
where $\tau$ is a positive definite diagonal matrix with its diagonal elements as $\text{diag}(\tau)=[\tau_{1},\tau_{2},\tau_{3}]$ and $\mu$, $\Lambda$ and $\Theta$ are positive definite matrices.
For stability analysis, we consider a Lyapunov function candidate as
$V=$ $\frac{1}{2}s^{T}s+\frac{1}{2}\tilde{d}^{T}\tilde{d}+\frac{1}{2}\tilde{k}^{T}\tilde{k}$
$\implies\dot{V}=$ $s^{T}\left(J^{-1}\tilde{d}-\Theta s\right)+\tilde{d}^{T}\dot{d}-\tilde{d}^{T}\Lambda J^{-1}\tilde{d}-\tilde{d}^{T}\hat{k}\text{sgn}(\tilde{d})$
$-\tilde{d}^{T}J^{-1}s-\tilde{k}^{T}\dot{\hat{k}}$
$\leqslant$ $-s^{T}\Theta s+|\tilde{d}^{i}|k_{i}-\tilde{d}^{T}\Lambda J^{-1}\tilde{d}-|\tilde{d}^{i}|\hat{k}_{i}-\tilde{k}^{T}\dot{\hat{k}}$
$=$ $-s^{T}\Theta s+|\tilde{d}^{i}|\tilde{k_{i}}-\tilde{d}^{T}\Lambda J^{-1}\tilde{d}-\tilde{k}^{T}\dot{\hat{k}}$
$\leqslant$ $-s^{T}\Theta s+\frac{1}{2}\tilde{d}^{T}\tilde{d}+\frac{1}{2}\tilde{k}^{T}\tilde{k}-\tilde{d}^{T}\Lambda J^{-1}\tilde{d}$
$+\tilde{k}^{T}(\tau\hat{k}-\mu|s|)$
$=$ $-s^{T}\Theta s-\tilde{d}^{T}\left(\Lambda J^{-1}-\frac{1}{2}I_{3}\right)\tilde{d}+\frac{1}{2}\tilde{k}^{T}\tilde{k}$
$+\hat{k}^{i}\tau_{i}\tilde{k}_{i}-|s|\mu\tilde{k}.$ (36)
Using the fact that
$\tilde{k}^{T}\mu|s|\leqslant\frac{\lambda_{\text{max}}(\mu)}{2}\left(\tilde{k}^{T}\tilde{k}+s^{T}s\right),$
(36) can be expressed as
$\dot{V}\leqslant$ $-s^{T}\left(\Theta-\frac{1}{2}\lambda_{\text{max}}(\mu)I_{3}\right)s-\tilde{d}^{T}\left(\Lambda J^{-1}-\frac{1}{2}I_{3}\right)\tilde{d}$
$+\frac{1}{2}\tilde{k}^{T}\tilde{k}+\hat{k}^{i}\tau_{i}\tilde{k}_{i}+\frac{\lambda_{\text{max}}(\mu)}{2}\tilde{k}^{T}\tilde{k}$
$=$ $-s^{T}\left(\Theta-\frac{1}{2}\lambda_{\text{max}}(\mu)I_{3}\right)s-\tilde{d}^{T}\left(\Lambda J^{-1}-\frac{1}{2}I_{3}\right)\tilde{d}$
$+\left(\frac{\lambda_{\text{max}}(\mu)+1}{2}\right)\tilde{k}^{T}\tilde{k}+\tau^{i}\left(k_{i}\hat{k}_{i}-\hat{k_{i}}^{2}\right).$ (37)
Consider the following inequality
$k_{i}\hat{k}_{i}-\hat{k}_{i}^{2}$ $=-\left(\frac{k_{i}}{\sqrt{2}}-\frac{\hat{k}_{i}}{\sqrt{2}}\right)^{2}-\frac{\hat{k}_{i}^{2}}{2}+\frac{k_{i}^{2}}{2},$
$\implies$ $\leqslant-\frac{\tilde{k}_{i}^{2}}{2}+\frac{k_{i}^{2}}{2}.$ (38)
Hence, by substituting the above (38) in equation (37)
$\dot{V}\leqslant$ $-s^{T}\left(\Theta-\frac{\lambda_{\text{max}}(\mu)}{2}I_{3}\right)s-\tilde{d}^{T}\left(\Lambda J^{-1}-\frac{1}{2}I_{3}\right)\tilde{d}$
$-\frac{1}{2}\tilde{k}^{T}\left(\tau-\lambda_{\text{max}}(\mu)I_{3}-I_{3}\right)\tilde{k}+\frac{1}{2}k^{T}\tau k$
$\implies\dot{V}\leqslant$ $-\Gamma V+\bar{\Delta},$ (39)
where $\Gamma=\text{min}\left(\lambda_{\text{min}}(\bar{\Theta}),\lambda_{\text{min}}(\bar{\Lambda}),\lambda_{\text{min}}(\frac{\tau_{0}}{2})\right)$, $\bar{\Theta}=\Theta-\frac{\lambda_{\text{max}}(\mu)}{2}I_{3}$, $\bar{\Lambda}=\Lambda J^{-1}-\frac{1}{2}I_{3}$ and $\bar{\Delta}=\lambda_{\text{max}}(\tau)||k||^{2}$. On similar lines to that of eq. (LABEL:pf_3) and (LABEL:lya_guub1), eq (39) can be expressed as
$\dot{V}\leqslant-\varrho V-(\Gamma-\varrho)V+\bar{\Delta}.$
For $\dot{V}(t)\leqslant 0$ for $V(0)>\mathcal{R}_{\text{sc}}$, there exists some $\mathcal{R}_{\text{sc}}=\frac{\bar{\Delta}}{\Gamma-\varrho}$ such that $V(t)$ reaches the closed ball of radius $\mathcal{R}_{\text{sc}}$ in finite time, say $t_{r_{\text{sc}}}$ and remain there $\forall t>t_{r_{\text{sc}}}$. $t_{r_{\text{sc}}}$ can be evaluated as
$t_{r_{\text{sc}}}$ $\leqslant\frac{1}{\Gamma}\text{ln}\left[\frac{V(0)-\frac{\bar{\Delta}}{\Gamma}}{\bar{\Delta}\left(\frac{1}{\Gamma-\varrho}-\frac{1}{\Gamma}\right)}\right].$
As the sliding variable component in the Lyapunov function, $\frac{1}{2}s^{2}\leqslant V$, the state trajectories are also contained within a closed ball of radius $\sqrt{\frac{2\bar{\Delta}}{\Gamma-\varrho}}$.
(a) Evolution of quaternions
(b) Evolution of angular velocities
(c) Sliding variable of FOITSM for rigid spacecraft dynamics
(d) Disturbance estimation error of adaptive ADO with FOITSMC
(e) Estimation of $k$
(f) Input for FOITSMC with adaptive ADO
Figure 5: 5a, 5b and 5c show the global boundedness of quaternions, $\Omega$, and sliding variable, respectively, in the presence of matched disturbances. 5d shows the global boundedness of disturbance estimation error. The continuous control using adaptive ADO-based FOITSMC and the estimated value of $k$ is shown in 5e and 5f, respectively.
The parameters used in the simulation are as follows: $\Theta=2I_{3}$, $\Lambda=50I_{3}$, $\mu=2I_{3}$, $\tau=5I_{3}$ and the inertia matrix,
$J=\begin{pmatrix}20&0&0.9\\ 0&17&0\\ 0.9&0&15\end{pmatrix}$
The simulation results shown in fig. 5a and 5b, utilising the ODE 4 (Runge-Kutta) solver of MATLAB R2021a with an assigned fixed step size of 0.001, depict that the state trajectories remain close to the origin, bounded within a specified limit. A continuous control input has been employed using an adaptive ADO-based FOITSMC algorithm, as shown in Fig. 5f, to ensure the global boundedness of error in the estimation of conditionally known disturbance and unknown gain $k$.
8 Conclusion and outlook
This paper presents an algorithm ensuring the Global Uniform Ultimate Boundedness (GUUB) of state trajectories in the presence of disturbances with unknown bounds. Additionally, a sliding manifold is devised that combines the attributes of Integral Terminal Sliding Mode Control (ITSMC) and nonsingular Terminal Sliding Mode Control (TSMC). The Lyapunov-based stability analysis incorporates the dynamics of the Adaptive Disturbance Observer (ADO). Consequently, the monotonically increasing nature of the adaptive gain is eliminated. For future work, exploring the effectiveness of the algorithm for dynamical systems that are not affine in control input, designing an observer-based controller for the same, modifying the algorithm to work in the realm of mismatched perturbations, and determining a more stringent bound on states and errors can be a potential research objective.
Disclosure statement
The authors have no conflict of interest to disclose.
Appendix
8.1 Proof of Theorem 2.3
Differentiating (6) once with respect to time
$\dot{s}=$ $\dot{x}_{n}+C_{n}|x_{n}|^{\alpha_{n}}\text{sgn}(x_{n})+C_{n-1}|x_{n-1}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})$
$+...+C_{1}|x_{1}|^{\alpha_{1}}\text{sgn}(x_{1}).$
Substituting the dynamics of $x_{n}$ from (5)
$\dot{s}$ $=f_{n}(\textbf{x},t)+b(\textbf{x},t)u+d(\textbf{x},t)+C_{n}|x_{n}|^{\alpha_{n}}\text{sgn}(x_{n})$
$+C_{n-1}|x_{n-1}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})+...+C_{1}|x_{1}|^{\alpha_{1}}\text{sgn}(x_{1}).$
Substituting $u$, $u_{eqv}$, and $u_{dis}$, respectively, from (17), (18) and (19),
$\dot{s}$ $=d(\textbf{x},t)+f_{\Delta}(\textbf{x},t)-(\eta+d_{max}+f_{max})\text{sgn}(s),$
$\leqslant-\eta\;\text{sgn}(s),$
which implies that $s$, if it was initially at zero, remains at zero for all time $t\geqslant 0$. Hence, the system evolves by following the dynamical equation (16) from the initial time $t=0$. This completes the proof.
8.2 Proof of Theorem 4.1
Consider a Lyapunov function candidate as
$V=$ $\frac{1}{2}s^{2}+\frac{1}{2}\tilde{d}^{2}+\frac{1}{2}\tilde{k}^{2},$
$\dot{V}=$ $s\left(\tilde{d}-\kappa s\right)+\tilde{d}\left(\dot{d}-\dot{\hat{d}}\right)+\tilde{k}(-\dot{\hat{k}}),$
$=$ $-\kappa s^{2}+s\tilde{d}+\tilde{d}\left(\dot{d}-\lambda\tilde{d}-\hat{k}\text{sgn}(\tilde{d})-s\right)$
$+\tilde{k}(-\dot{\hat{k}}),$
$\leqslant$ $-\kappa s^{2}-\lambda\tilde{d}^{2}+\frac{1}{2}\tilde{k}^{2}+\frac{1}{2}\tilde{d}^{2}+\tilde{k}\left(\tau\hat{k}-\mu|s|\right),$
$=$ $-\kappa s^{2}-\left(\lambda-\frac{1}{2}\right)\tilde{d}^{2}+\frac{1}{2}\tilde{k}^{2}+\tilde{k}\left(\tau\hat{k}-\mu|s|\right),$
$\leqslant$ $-\kappa s^{2}-\left(\lambda-\frac{1}{2}\right)\tilde{d}^{2}+\frac{1}{2}\tilde{k}^{2}+\tau\tilde{k}\hat{k}+\frac{1}{2}\mu\tilde{k}^{2}+\frac{1}{2}\mu s^{2},$
$=$ $-\left(\kappa-\frac{1}{2}\mu\right)s^{2}-\left(\lambda-\frac{1}{2}\right)\tilde{d}^{2}+\frac{1}{2}(\mu+1)\tilde{k}^{2}+\tau\tilde{k}\hat{k},$
$=$ $-\left(\kappa-\frac{1}{2}\mu\right)s^{2}-\left(\lambda-\frac{1}{2}\right)\tilde{d}^{2}+\frac{1}{2}(\mu+1)\tilde{k}^{2}$
$+\tau\left(k\hat{k}-\hat{k}^{2}\right).$ (40)
Using the fact that $k\hat{k}\geqslant 0$ for all time and,
$k\hat{k}-\hat{k}^{2}$ $=-\left(\frac{k}{\sqrt{2}}-\frac{\hat{k}}{\sqrt{2}}\right)^{2}-\frac{\hat{k}^{2}}{2}+\frac{k^{2}}{2},$
$\implies$ $\leqslant-\frac{1}{2}\tilde{k}^{2}+\frac{k^{2}}{2}.$ (41)
Using (41), (40) becomes
$\dot{V}\leqslant$ $-\left(\kappa-\frac{1}{2}\mu\right)s^{2}-\left(\lambda-\frac{1}{2}\right)\tilde{d}^{2}$
$-\frac{1}{2}\left(\tau-\mu-1\right)\tilde{k}^{2}+\frac{1}{2}\tau k^{2},$
$\implies\dot{V}\leqslant$ $-\gamma V+\bar{\delta},$
where $\gamma=\text{min}\left\{\bar{\kappa},\bar{\lambda},\frac{\tau_{0}}{2}\right\}>0$, $\bar{\kappa}=\kappa-\frac{1}{2}\mu$, $\bar{\lambda}=\lambda-\frac{1}{2}$, $\tau_{0}=\left(\tau-\mu-1\right)$ and $\bar{\delta}=\frac{1}{2}\tau k^{2}$.
For $0<\theta<\gamma$, the above equation can be rewritten as,
$\dot{V}$ $=-\theta V-\left(\gamma-\theta\right)V+\bar{\delta}.$
For some $\mathcal{R}=\frac{\bar{\delta}}{\gamma-\theta}$, $\dot{V}(t)<0$ for $V(0)>\mathcal{R}$. Hence $V(t)$ reaches $\mathcal{R}$ in finite time, say $t_{r}$ and remains there in the closed ball of radius $\mathcal{R},\;\forall t>t_{r}$, where $t_{r}$ can be evaluated as given below
$t_{r}$ $\leqslant\frac{1}{\gamma}\text{ln}\left[\frac{V(0)-\frac{\bar{\delta}}{\gamma}}{\bar{\delta}\left(\frac{1}{\gamma-\theta}-\frac{1}{\gamma}\right)}\right].$ (42)
As the term containing sliding variable $\frac{1}{2}s^{2}\leqslant V$, the state trajectories are also confined within the closed ball of radius $\sqrt{\frac{2\bar{\delta}}{\gamma-\theta}}$. This ends the proof.